Log2X + Logx2 = 2, find x


1
Expert's answer
2021-08-30T16:23:40-0400

"\log_2x + log_x 2 =2
\\= log_2x + \frac{1}{\log_2x}= 2
\\\text{Let $\log_2x=y$}
\\y + \frac{1}{y}=2
\\=y^2 +1=2y
\\=y^2 -2y+1=0
\\(y-1)^2=0
\\y=1
\\\text{Recall that $log_2x=1$}
\\\implies x =2"